OpenStudy (istim):

Determine the work done by each force in newtons for an obect moving along the vector s, in metres. F=[67.8,3.9],s=[4.7,3.2]

OpenStudy (istim):

\[\cos^-1(44.346/(\sqrt{71.7}\sqrt{7.9}))=\theta\]

OpenStudy (istim):

That's as far as I got, but then I got an error on my calculator. Error 2.

OpenStudy (istim):

I guess you rather me show my work completely, no?

OpenStudy (anonymous):

how far is the object moving?

OpenStudy (istim):

I've given the data completely.

OpenStudy (istim):

The question here is not that of physics, but of cartesian vectors.

OpenStudy (anonymous):

ok. well if s is the displacement vector, then you have, \[W=F*s=<67.8,3.9>*<4.7,3.2>\]This gives: \[W=318.66+12.48=331.14J\]

OpenStudy (istim):

But I'm suppose to do it the long, complex, Cartesian way!

OpenStudy (istim):

I had to use these formulas- \[F \times s= F \times s \cos \Theta\]

OpenStudy (istim):

And: \[W=F \times s \cos #\]

OpenStudy (anonymous):

Are you guys using dot products? because W=(F)(s)cosa is equivalent to my calculation.

OpenStudy (istim):

Yeah. We're doing dot products!

OpenStudy (anonymous):

I don't think I'm following what the problem is. Just dot all the things.

OpenStudy (istim):

Complete steps would be appreciated, but I understand that it's long and tedious. Maybe I should give a complete info on my course next time.

OpenStudy (istim):

"Dot all the things"?

OpenStudy (istim):

Maybe I should show my complete steps, and you could tell me what I did wrong?

OpenStudy (anonymous):

Ok, well that is how its done. when you are given two vectors in component form, you just multiply the components and add the results like I did above. The version with the cos in it is only useful if we are given the angle between the vectors or if they want the angle between the vectors.

OpenStudy (istim):

SHOOT. AFK. Brb

OpenStudy (anonymous):

k

OpenStudy (istim):

in 10 min or so.

OpenStudy (istim):

Go do important things

OpenStudy (anonymous):

Well, the original question gave us a position vector with what I think is the magnitude of its components. And then a force vector with its component's magnitudes. By finding the force vector's component in the position vector's resultant direction, we can find work. i.e., dotting all the things

OpenStudy (anonymous):

The only way this question makes sense is if s is the displacement vector, not the position vector. But otherwise, yeah, just dot product them using the given components.

OpenStudy (anonymous):

My bad, yeah, displacement vector.

OpenStudy (istim):

Ok. Thank you guys.

OpenStudy (anonymous):

The only way this question makes sense is if s is the displacement vector, not the position vector. But otherwise, yeah, just dot product them using the given components.

OpenStudy (istim):

Help appreciated.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

np